OpenStudy (nicholausblackmon):

I am holding a pop quiz. Whoever answers correctly first gets a fan and a medal!

OpenStudy (nicholausblackmon):

when five people join i post answer

OpenStudy (anonymous):

hellooooooooo were is the question

OpenStudy (nicholausblackmon):

5(-3x - 2) - (x - 3) = -4(4x + 5) + 13

OpenStudy (nicholausblackmon):

find "x"

OpenStudy (nicholausblackmon):

@1568

OpenStudy (anonymous):

i don't know what its called but -7=-7

OpenStudy (nicholausblackmon):

Nope!

OpenStudy (nicholausblackmon):

@1497

OpenStudy (mendicant_bias):

\[5(-3x-2)-(x-3)=-4(4x+5)=13;\] \[-15x-10-x+3=-16x-20+13.\]

OpenStudy (mendicant_bias):

The whole thing cancels out. You get zero equals zero; you cannot solve this equality for a given value of x, because combining like terms just results in\[0=0.\]

OpenStudy (anonymous):

Isnt x all numbers?

OpenStudy (nicholausblackmon):

You win!

OpenStudy (mendicant_bias):

(Who, lol.)

OpenStudy (anonymous):

who wins

OpenStudy (mendicant_bias):

Marie is also entirely right, what she said is true, too. x can be any value at all. It's a trivial solution, but it's still entirely true.

OpenStudy (anonymous):

Yup